OpenStudy (anonymous):

simplify the expression. 6 to the square root of 8 * 7 to the square root of 2

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Welcome!

OpenStudy (anonymous):

\[42*\sqrt{8*2}\]=168

OpenStudy (anonymous):

\[6\sqrt{8} * 7\sqrt{2}\]

OpenStudy (anonymous):

hurry up

OpenStudy (anonymous):

6* 8 * 7*2^0.5 0 ... 6 * 2 *2^0.5 * 7*2^0.5 0 ... 12 * 7 *2^0.5 * 2^0.5 ... 12 * 7 * 2 .... 168 NB 2^0.5 means the sqrt of 2

OpenStudy (anonymous):

i dont get it?

OpenStudy (anonymous):

Which parrt?

OpenStudy (anonymous):

all of it haha

OpenStudy (anonymous):

Give me 5 minutes and I will post a detailed jpeg oh and stop telling kind teachers to hurry up :)

OpenStudy (anonymous):

sorry.

OpenStudy (anonymous):

OK I will be back soon

OpenStudy (anonymous):

k hurry up lol

OpenStudy (anonymous):

so what is it?

OpenStudy (anonymous):

whats taking so long..

OpenStudy (anonymous):

Word doc not jpeg

OpenStudy (anonymous):

Any better?

OpenStudy (anonymous):

Hurry up

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

\[\sqrt{20} /\sqrt{5}\]

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

Root 20 is root 4 x root 5 Root 4 is 2 So the fraction is now 2 root 5/root 5 which is ....2

OpenStudy (anonymous):

Check on your calculator.

OpenStudy (anonymous):

Come on now I am getting tired of waiting :D

OpenStudy (anonymous):

dont get lipy with me

OpenStudy (anonymous):

Medals usually keep me quiet.

OpenStudy (anonymous):

u dont deserve em u suck

OpenStudy (anonymous):

You certainly know how to get people to help you

OpenStudy (anonymous):

sorry its 4 am...

OpenStudy (anonymous):

Where

OpenStudy (anonymous):

minnesota

OpenStudy (anonymous):

where r u?

OpenStudy (anonymous):

Ok I forgive you, but did you understand the last one?

OpenStudy (anonymous):

France

OpenStudy (anonymous):

cool haha whats ur name?

OpenStudy (anonymous):

Gianfranco

OpenStudy (anonymous):

is that ur real name?

OpenStudy (anonymous):

Yes it is italian but it is lunchtime in France and that is very important to our cultue so I will have to go now. Anymore problems with maths I will be glad to help.

OpenStudy (anonymous):

alright thanks

OpenStudy (anonymous):

Ciao bon journee!

OpenStudy (anonymous):

whats that?

OpenStudy (anonymous):

So long have a good day!

OpenStudy (anonymous):

oh haha u tooo

OpenStudy (anonymous):

6sqrt8*7sqrt2 we can say sqrt8= sqrt2x2x2=2sqrt2 then we have 6*2sqrt2*7sqrt2 84*2 168 ans